The cube shown here has volume 1000 cm3. If its edges are all doubled in length its new volume will be
Answer by stanbon(75887)

The cube shown here has volume 1000 cm3. If its edges are all doubled in length its new volume will be
----
Original edge length = 1000^(1/3) = 10 cm
----
New edge length = 20
----
New volume = 20^3 = 8000 cm^3
=====================
Cheers,
Stan H.